# Generating, Reconstructing, and Representing Discrete and Continuous Data: Generalized Diffusion with Learnable Encoding-Decoding

### v1

@@ 6 Conclusion
In this work, we generalized the diffusion model to introduce DiLED, an innovative generative framework designed to seamlessly integrate the three core functionalities for generative models: generation, reconstruction and representation. To achieve the goad, we incorporated parameterized encoder and decoder transformations into the conventional diffusion process, redefining both the forward and reverse process. Subsequently, we derived an end-to-end training objective from the data likelihood, which encompasses three components, each addressing one of the core functionalities. Experimental results across image, text, and protein sequence data demonstrate that DiLED consistently outperforms strong baselines across all three core functionalities. Building on the versatility of our approach, which accommodates a range of data types via tailored encoder/decoder configurations, future investigations could extend its applicability across diverse domains.
@@ A.4 Derivation of our Final loss
$\mathcal{L}^{\text{final}}=\mathbb{E}_{q,{\bm{\epsilon}}}\left[\sum_{t=3}^{T}\gamma_{t}\left(w\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\left({\mathbf{x}}_{t},t\right)\right\|^{2}+\frac{1}{\gamma_{t}T}\mathcal{L}_{\text{rec}}\right)+\gamma_{1}\left(w\left(\rho\|\mathcal{E}_{\lambda}({\mathbf{x}}_{0})-{\bm{\mu}}_{\theta}({\mathbf{x}}_{2},1)\|^{2}\right)+\frac{1}{\gamma_{1}T}\mathcal{L}_{\text{rec}}\right)\right].$ (33)
@@ Appendix B Algorithm
Below shows the complete training algorithm of DiLED.
@@ C.1.1 Setup
In our study, we aim to comprehensively assess the holistic performance of DiLED by evaluating it across three foundational capabilities: generation, reconstruction, and latent vector arithmetic. To achieve this, we focus on four specific tasks: generation, reconstruction, interpolation, and text style transfer using latent vector arithmetic. Each of these tasks is selected for the following reasons:
@@
  • Fluent Generation: The ultimate test of our model’s effectiveness lies in the quality of its output. For DiLED, the fluency and coherence of the generated text are crucial indicators of its robust generative capabilities. This is demonstrated in generation tasks, text style transfer, and interpolation.
By evaluating DiLED on these four tasks, we can effectively reflect and analyze its performance in terms of the three core functionalities.
@@ C.1.2 Detailed Results
Reconstruction Generation Style Transfer (Arithmetic)
@@
C.1.3 Training Curves
Comparing to other methods, our training is more stable and converges faster as illustrated in Fig. 3.
@@ C.2.2 Detailed Experimental Results
The results are shown in Table 6. Note that we draw Figure 4 according to the generation, reconstruction performances, as well as the interpolation results with $\alpha=0.4$.
@@
FFHQ 128 LDM 67.78 30.43 12.90 4.87 21.29 75.13
NVAE 41.26 5.73 6.28 9.34
StyleGAN-XL 24.72 34.09 46.82 59.64
@@
DiLED 18.41 14.38 12.26 11.50 8.17 5.48 12.57 9.80 6.66 19.11 18.36 16.98
CelebA 64 LDM 41.87 31.40 25.80 9.58 9.95 28.78
NVAE 23.11 1.55 3.03 8.49
@@
DiLED 11.01 8.03 6.35 6.21 4.77 3.49 7.26 6.31 5.39 13.20 11.83 10.58
Table 6: Overall performance comparison in various tasks. We use FID and rFID to evaluate the performances of the generation and reconstruction, respectively. For the interpolation tasks, we test two settings $\alpha=0.2$ and $\alpha=0.4$, where FID is also used to evaluate the performance.
@@ C.2.4 Image Reconstruction
For a comparative analysis, we present reconstructed images from various models, each utilizing a distinct total step $T$ in the decoder. The results for FFHQ128 and CelebA64 are depicted in Figure 15 and Figure 16, respectively. From the figures, it’s clear that the VAE in LDM is effective at reconstruction. Our model DiLED also produces strong results.
@@ C.2.5 Image Representation
Class # Positives DiffAE DiLED
@@ C.3.1 Setup
In addition to ReLSO, an autoencoder-based model introduced above, we trained a vanilla Variational Autoencoder(Kingma & Welling, 2014) consists of the same transformer as encoder and convolutional layers as decoder. The dimension of the latent space of ReLSO and VAE are all set to 30. Furthermore, we have conducted comparisons with latest advancements in protein sequence diffusion models, namely NOS (Gruver et al., 2023). NOS proposes a Discrete Diffusion Model that based on BERT and utilized [MASK] as noise, and a Gaussian Diffusion Model that jointly learned a word embedding function and used BERT as the backbone of the diffusion model. While ReLSO, VAE and DiLED utilized a regressor of linear layers and ReLU, the regressor of NOS required an additional BertPooler to aggregate a sequence of latent embeddings. We have conducted experiments on Protein Representation with DiLED, ReLSO, VAE and NOS; Protein Optimization with DiLED, ReLSO and VAE; Protein Reconstruction with DiLED, ReLSO and VAE; and Protein Generation with DiLED, VAE and NOS.
@@
DiLED ($w_{\text{protein}}$=0.5) 0.231 0.346 0.840 0.477 1.016
DiLED ($w_{\text{protein}}$=5) 0.211 0.331 0.844 0.472 1.143
@@
DiLED ($w_{\text{protein}}$=0.5) 0.496 0.464 0.854 0.753 0.095
DiLED ($w_{\text{protein}}$=0.5) 0.488 0.440 0.848 0.734 0.094
@@ C.3.2 Protein Optimization
DiLED ($w_{\text{protein}}$=0.5) 2.003 2.000 12.615 6 29.219
DiLED ($w_{\text{protein}}$=5) 2.000 2.000 12.256 4 29.452
@@
DiLED ($w_{\text{protein}}$=0.5) 4.001 4.000 12.836 1 23.488
DiLED ($w_{\text{protein}}$=5) 4.001 4.000 1.076 1 23.480
@@
DiLED ($w_{\text{protein}}$=0.5) 1.000 1.000 13.35 6 29.352
@@
DiLED ($w_{\text{protein}}$=0.5) 1.501 1.500 13.07 6 29.294
@@
DiLED ($w_{\text{protein}}$=0.5) 2.003 2.000 12.61 6 29.219
@@
DiLED ($w_{\text{protein}}$=0.5) 2.504 2.500 12.30 6 28.736
@@
DiLED ($w_{\text{protein}}$=0.5) 3.001 3.000 3.41 1 23.478
@@
DiLED ($w_{\text{protein}}$=0.5) 4.000 4.000 12.83 1 23.488
@@
DiLED ($w_{\text{protein}}$=0.5) 1.003 1.000 13.067 6 29.200
DiLED ($w_{\text{protein}}$=5) 1.004 1.000 12.486 5 29.490
@@
DiLED ($w_{\text{protein}}$=0.5) 3.007 3.001 4.173 1 23.481
DiLED ($w_{\text{protein}}$=5) 3.003 3.002 0.715 1 23.475
@@ C.3.3 Protein Reconstruction
Reconstruction is evaluated with Cross Entropy to compare input and reconstructed protein sequences in the test set. The results are presented in the last column of Table 8. From Table 8, DiLED achieves comparable reconstruction ability compared to the baselines. Note that since NOS is a transformer-based model with word embedding that could not perform reconstruction in the same settings as other models.
@@ C.3.4 Conditional Protein Generation
For conditional generation with VAE and DiLED, latent representations are sampled and optimized with LatentOps algorithm with target fitness of 1 and 3 for Gifford and GFP dataset respectively. For NOS, the latents are updated through out each step of the diffusion process with Langevin Dynamics using gradients of the regressor. The quantitative evaluation are presentsed in 12. Although NOS models achieved better NLL, the generated proteins are have lower fitness values than other methods.
@@ D.2 Latent Score-based Generative Model
The Latent Score-based Generative Model (LSGM) enhances generative capabilities by learning Score-based Generative Models (SGM) within the latent space of a Variational Autoencoder (VAE). We will delineate the detailed differences between LSGM and DiLED in the following sections.
@@
LSGM: Their method can be regarded as a Variational Autoencoder (VAE) with a Score-based Generative Model (SGM) as the prior. Consequently, their objective is derived from Eq. (37). However, they uniquely decompose the KL term into two components, and each component is approximated based on certain assumptions. This approach results in a fundamentally different final training objective to that of DiLED.

### v2

@@ 6 Conclusion
This work generalized the diffusion model to introduce EDDPMs, an innovative generative framework designed to seamlessly integrate the three core functionalities for generative models: generation, reconstruction and representation. To this end, we incorporated parameterized encoder and decoder transformations into the conventional diffusion process. We derived an end-to-end training objective from the data likelihood inspired from the widely-used DDPM training framework. Experimental results across image, text, and protein sequence data demonstrate that EDDPMs consistently outperform strong baselines across all three core functionalities. Building a compact high-quality representation space with the versatile utilities (generation, reconstruction, encoding, interpolation, editing) can be significant for building new foundation models, such as world models (Hu & Shu, 2023; LeCun, 2022; Assran et al., 2023), for more robust machine reasoning (Hao et al., 2023; Wong et al., 2023).
Limitations
In this work, EDDPMs have not been verified on other data modalities such as videos, audios, and time series. It would be interesting to see how EDDPMs will perform on all the diverse forms of data, and even multi-modal data (such as text-image). In addition, we have only parameterized the first diffusion step with learnable encoding/decoding. A single latent vector representation could be limited to capture relevant information of data for different tasks. We would like to investigate our approach with more encoding/decoding-based diffusion steps for hierarchical latent representations.
Impact Statement
There are potential societal consequences associated with diffusion modeling and deep generative models in general regarding content creation, privacy, and others.
@@ A.4 Derivation of our Final loss
$\mathcal{L}^{\text{final}}=\mathbb{E}_{q,{\bm{\epsilon}}}\left[\sum_{t=3}^{T}\underbrace{\gamma_{t}\left(w\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\left({\mathbf{x}}_{t},t\right)\right\|^{2}+\frac{1}{\gamma_{t}T}\mathcal{L}_{\text{rec}}\right)}_{\mathcal{L}^{\text{final}}_{t-1}}+\underbrace{\gamma_{1}\left(w\left(\rho\|\mathcal{E}_{\lambda}({\mathbf{x}}_{0})-{\bm{\mu}}_{\theta}({\mathbf{x}}_{2},1)\|^{2}\right)+\frac{1}{\gamma_{1}T}\mathcal{L}_{\text{rec}}\right)}_{\mathcal{L}^{\text{final}}_{1}}\right].$ (33)
@@ Appendix B Algorithm
Below shows the complete training algorithm of EDDPMs.
@@ C.1.1 Setup
In our study, we aim to comprehensively assess the holistic performance of EDDPMs by evaluating it across three foundational capabilities: generation, reconstruction, and latent vector arithmetic. To achieve this, we focus on four specific tasks: generation, reconstruction, interpolation, and text style transfer using latent vector arithmetic. Each of these tasks is selected for the following reasons:
@@
  • Fluent Generation: The ultimate test of our model’s effectiveness lies in the quality of its output. For EDDPM, the fluency and coherence of the generated text are crucial indicators of its robust generative capabilities. This is demonstrated in generation tasks, text style transfer, and interpolation.
By evaluating EDDPM on these four tasks, we can effectively reflect and analyze its performance in terms of the three core functionalities.
@@ C.1.2 Detailed Results
Reconstruction Generation ——- Style Transfer (Arithmetic) ——-
@@ C.2.2 Detailed Experimental Results
The results are shown in Table 7. Note that we draw Figure 4 according to the generation, reconstruction performances, and interpolation results with $\alpha=0.4$. As for the results of CM-CD, the reported FID for generation in the original paper (Song et al., 2023) is 5.22, whereas we obtained an FID of 7.01. This discrepancy arises due to two main reasons: (1) We calculate the reference statistics (mean and variance) using the original images, resulting in slightly different statistics compared to those provided here [^4] https://github.com/openai/guided-diffusion/tree/main/evaluations. (2) The original model was trained on images with a resolution of 256x256. Consequently, we need to convert the generated 256x256 images to 128x128 for evaluation, which may introduce differences in FID. When using the pre-calculated statistics and evaluating the model at a resolution of 256x256, we obtain an FID of 5.82, which is close to the 5.22 reported in the original paper. Using the newly calculated statistics, the FID is 6.56. Upon converting the images to 128x128, the FID becomes 7.01.
@@
FFHQ 128 LDM 67.78 30.43 12.90 ——- 4.87 ——- ——- 21.29 ——- ——- 75.13 ——-
NVAE ——- 41.26 ——- ——- 5.73 ——- ——- 6.28 ——- ——- 9.34 ——-
StyleGAN-XL ——- 24.72 ——- ——- 34.09 ——- ——- 46.82 ——- ——- 59.64 ——-
@@
EDDPM 18.41 14.38 12.26 11.50 8.17 5.48 12.57 9.80 6.66 19.11 18.36 16.98
CelebA 64 LDM 41.87 31.40 25.80 ——- 9.58 ——- ——- 9.95 ——- ——- 28.78 ——-
NVAE ——- 23.11 ——- ——- 1.55 ——- ——- 3.03 ——- ——- 8.49 ——-
@@
EDDPM 11.01 8.03 6.35 6.21 4.77 3.49 7.26 6.31 5.39 13.20 11.83 10.58
Table 7: Overall performance comparison in various tasks. We use FID and rFID to evaluate the performances of the generation and reconstruction, respectively. For the interpolation tasks, we test two settings $\alpha=0.2$ and $\alpha=0.4$, where FID is also used to evaluate the performance. We highlight the best performance within the performances of methods that need to set up the parameter $T$ with bold. The performances of the methods with only one step such as NVAE, StyleGAN-XL, and CM-CD are underlined if their performances are better than the best performances of other methods with $T=50$.
@@ C.2.4 Image Reconstruction
For a comparative analysis, we present reconstructed images from various models, each utilizing a distinct total step $T$ in the decoder. The results for FFHQ128 and CelebA64 are depicted in Figure 15 and Figure 16, respectively. From the figures, it’s clear that the VAE in LDM is effective at reconstruction. Our model EDDPM also produces strong results.
@@ C.2.5 Image Representation
Class # Positives DiffAE EDDPM
@@ C.3.1 Setup
In addition to ReLSO, an autoencoder-based model introduced above, we trained a vanilla Variational Autoencoder(Kingma & Welling, 2014) consists of the same transformer as encoder and convolutional layers as decoder. The dimension of the latent space of ReLSO and VAE are all set to 30. Furthermore, we have conducted comparisons with latest advancements in protein sequence diffusion models, namely NOS (Gruver et al., 2023). NOS proposes a Discrete Diffusion Model that based on BERT and utilized [MASK] as noise, and a Gaussian Diffusion Model that jointly learned a word embedding function and used BERT as the backbone of the diffusion model. While ReLSO, VAE and EDDPM utilized a regressor of linear layers and ReLU, the regressor of NOS required an additional BertPooler to aggregate a sequence of latent embeddings. We have conducted experiments on Protein Representation with EDDPM, ReLSO, VAE and NOS; Protein Optimization with EDDPM, ReLSO and VAE; Protein Reconstruction with EDDPM, ReLSO and VAE; and Protein Generation with EDDPM, VAE and NOS.
@@
EDDPM ($w_{\text{protein}}$=0.5) 0.231 0.346 0.840 0.477 1.016
EDDPM ($w_{\text{protein}}$=5) 0.211 0.331 0.844 0.472 1.143
@@
EDDPM ($w_{\text{protein}}$=0.5) 0.496 0.464 0.854 0.753 0.095
EDDPM ($w_{\text{protein}}$=0.5) 0.488 0.440 0.848 0.734 0.094
@@ C.3.2 Protein Optimization
EDDPM ($w_{\text{protein}}$=0.5) 2.003 2.000 12.615 6 29.219
EDDPM ($w_{\text{protein}}$=5) 2.000 2.000 12.256 4 29.452
@@
EDDPM ($w_{\text{protein}}$=0.5) 4.001 4.000 12.836 1 23.488
EDDPM ($w_{\text{protein}}$=5) 4.001 4.000 1.076 1 23.480
@@
EDDPM ($w_{\text{protein}}$=0.5) 1.000 1.000 13.35 6 29.352
@@
EDDPM ($w_{\text{protein}}$=0.5) 1.501 1.500 13.07 6 29.294
@@
EDDPM ($w_{\text{protein}}$=0.5) 2.003 2.000 12.61 6 29.219
@@
EDDPM ($w_{\text{protein}}$=0.5) 2.504 2.500 12.30 6 28.736
@@
EDDPM ($w_{\text{protein}}$=0.5) 3.001 3.000 3.41 1 23.478
@@
EDDPM ($w_{\text{protein}}$=0.5) 4.000 4.000 12.83 1 23.488
@@
EDDPM ($w_{\text{protein}}$=0.5) 1.003 1.000 13.067 6 29.200
EDDPM ($w_{\text{protein}}$=5) 1.004 1.000 12.486 5 29.490
@@
EDDPM ($w_{\text{protein}}$=0.5) 3.007 3.001 4.173 1 23.481
EDDPM ($w_{\text{protein}}$=5) 3.003 3.002 0.715 1 23.475
@@ C.3.3 Protein Reconstruction
Reconstruction is evaluated with Cross Entropy to compare input and reconstructed protein sequences in the test set. The results are presented in the last column of Table 9. From Table 9, EDDPM achieves comparable reconstruction ability compared to the baselines. Note that since NOS is a transformer-based model with word embedding that could not perform reconstruction in the same settings as other models.
@@ C.3.4 Conditional Protein Generation
For conditional generation with VAE and EDDPM, latent representations are sampled and optimized with LatentOps algorithm with target fitness of 1 and 3 for Gifford and GFP dataset respectively. For NOS, the latents are updated through out each step of the diffusion process with Langevin Dynamics using gradients of the regressor. The quantitative evaluation are presentsed in 13. Although NOS models achieved better NLL, the generated proteins are have lower fitness values than other methods.
C.3.5 Influence of weight $w$
In the objective (Eq.33), the weight $w$ aims to balance the generation and reconstruction. As mentioned in §3.1, we adopt $w=8$ for text and $w=1$ for both image and protein. The choice of $w$ is primarily influenced by the different loss functions used for the reconstruction term $L_{Rec}$ (Eq.4) across various data types. For example, cross-entropy is used for text, while mean squared error (MSE) is used for images. To ensure that the different loss terms have similar scales, we selected $w$ values that roughly balanced their magnitudes. The $w$ values are then fixed during training. As a comparison, beta-VAE on text typically requires careful scheduling/annealing of $\beta$ values during training (Higgins et al., 2017b; Li et al., 2020), making the training more difficult and unstable.
C.3.6 Posterior Collapse
EDDPM can alleviate the posterior collapse issue on text sequences thanks to the improved formulation and training:
  • As discussed in §3.3, EDDPMs can be viewed as a VAE with a jointly learned diffusion model prior (instead of a standard Gaussian prior in vanilla VAEs). More formally, the EDDPMs’ objective in Eq.7 includes the $L_{\text{align}}$ term that explicitly encourages alignment between the encoder’s posterior distribution $q_{\lambda}(\mathbf{x}_{1}|\mathbf{x}_{0})$ and the diffusion model’s prior $p_{\theta}(\mathbf{x}_{1}|\mathbf{x}_{2})$. This learned prior is more flexible and helps avoid the unreasonable regularization posed by standard Gaussian prior that causes posterior collapse.
  • EDDPMs derive the training objective from the well-established DDPM framework, enabling stable and effective joint training of the encoder, decoder, and diffusion components. For example, as mentioned above, EDDPMs are robust to the balancing weight $w$. This is in contrast to the instability and sensitivity to hyperparameters in text-VAE training. To alleviate posterior collapse, training text-VAEs often requires tricks like beta-annealing (Bowman et al., 2016), free bits (Kingma et al., 2017), cyclic annealing schedule (Fu et al., 2019), and so on. This can be observed through the training loss curves in Figure 3.
Therefore, the posterior collapse issue is not obvious in EDDPMs.
@@ Appendix D Contrasting with Similar Works
In §3.3, we provide a high-level comparison of EDDPMs with many related models, highlighting the connections and differences between our approach and various existing methods. In §5, we present additional discussion of related works, focusing on recent efforts to combine VAEs, GANs, and diffusion models. We further mention the limitations of these approaches and how EDDPMs aim to overcome them. In this section, we would like to offer an in-depth discussion of EDDPMs with two closely related methods: Latent Diffusion Models and Latent Score-based Generative Models.
@@ D.2 Latent Score-based Generative Model
The Latent Score-based Generative Model (LSGM) enhances generative capabilities by learning Score-based Generative Models (SGM) within the latent space of a Variational Autoencoder (VAE). We will delineate the detailed differences between LSGM and EDDPMs in the following sections.
@@
LSGM: Their method can be regarded as a Variational Autoencoder (VAE) with a Score-based Generative Model (SGM) as the prior. Consequently, their objective is derived from Eq. (37). However, they uniquely decompose the KL term into two components, and each component is approximated based on certain assumptions. This approach results in a fundamentally different final training objective to that of EDDPMs.